# Statistical features of quantum chaos using the Krylov operator complexity

### v2

@@ Abstract
Recently the variance of Lanzcos coefficients in the Krylov space of an initial operator is recognized as an important indicator of quantum chaos. In this paper, we generate samples of random initial operators from given probability distribtions (GOE, GUE and the uniform distribution) and do statistics on the variance of Lanczos coefficients over these initial operators. Using the Sinai billiard with an integrability-breaking term, we propose two statistical quantities that have distinct behaviors as the system changes from nonchaotic to chaotic. One is the average correlation matrix $\langle x_{i}x_{j}\rangle$ of Lanczos coefficients, which exhibits different patterns in the nonchaotic and the chaotic regime. The other one is the resulting distribution of the variance of Lanzcos coefficients. In the nonchaotic case, the resulting distributions are almost overlapping together. In the chaotic case, they split into two well-separated groups. Furthermore, the resulting statistics of these two quantities are the Wishart distribution and the (rescaled) chi-square distribution respectively, which are independent of the distributions of initial operators and become the normal distribution in the case of large matrix size.
@@ 1 Introduction
The physics of complexity is important for understanding many different phenomena, ranging from condensed matter physics to black hole physics. Especially, the operator complexity captures the spread of an operator under time evolution and is an important indicator in the analysis of chaos. Recently the spread of an operator is explored in the Krylov space Parker:2018yvk , where the dynamics depends on the expansion coefficients of the operator in the Krylov basis and on the Lanczos coefficients generated in the process of constructing this Krylov basis. The Krylov complexity (K-complexity) Parker:2018yvk constructed from the expansion coefficients is proposed as an upper bound on the operator complexity, including the famous bound $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena:2015waa of the growth of the out-of-time-order correlation function (OTOC) 1969JETP28:1200L .
@@
Then it leads to an interesting connection between K-complexity and Random Matrix Theory (RMT). By adding an integrability-breaking term to the XXZ spin chain, the late-time saturation value of K-complexity is computed as the system changes from nonchaotic to chaotic behavior Rabinovici:2022beu . It approaches the behavior of similar values computed with Hamiltonian drawn from RMT ensembles Rabinovici:2022beu . Also, three different initial operators generating the Krylov space is randomly selected and compared there Rabinovici:2022beu . In Craps:2024suj , the K-complexity is also studied by choosing multiple initial operators.
@@
Along this line, an important question arises: what would happen for the Lanczos coefficients if we do statistics of initial operators over a given probability distribution. In this paper, we study this possibility by generating samples of initial operators from a given probability distribution (GOE, GUE and the uniform distribution) and analyzing the resulting Lanzcos coefficients. Interestingly, we find two statistical quantities that behave distinctively in the nonchaotic and the chaotic regime after averaging over the sample. One is the average correlation matrix $\langle x_{i}x_{j}\rangle$ of Lanczos coefficients. In the nonchaotic regime, it has a pattern that locally resembles the two-electron wave function of the Anderson localization Weinmann1995h2eOF , but this pattern disappears (or fading away by spreading out uniformly) in the chaotic regime. The other one is the resulting distribution of the variance $\sigma^{2}$. For samples of initial operators from different distributions, the resulting distributions of $\sigma^{2}$ are nearly the same in the nonchaotic case, but split into two well-separated groups in the chaotic case. Intuitively, the resulting plots of distributions of $\sigma^{2}$ are overlapped together in the nonchaotic case, and divide into two well-separated groups in the chaotic case. Furthermore, we find that the average correlation matrix $\langle x_{i}x_{j}\rangle$ satisfies the Wishart distribution and the resulting distribution of $\sigma^{2}$ satisfies a rescaled chi-square distribution. This is independent of the initial operator distributions and approaches the normal distribution as the size of the operator matrix increases.
@@
In detail, we consider the Sinai billiard, which is a typical billiard system originating from a quantization of classical chaos BERRY1981163 . The numerical computation is done in the energy representation and operators become matrices. Each time, a random initial operator $\mathcal{O}$ is sampled out by picking the corresponding random matrix from the following distributions: the GOE, the GUE, the uniform distribution of complex numbers (UCP), the uniform distribution of pure real numbers (URE) and of pure imaginary numbers (UIM). Then the Lanczos coefficients are computed in the Krylov space associated with this operator, which give the correlation matrix $x_{i}x_{j}$ and the variance $\sigma^{2}$ of Lanczos coefficients. Repeating this enough times, the average correlation matrix $\langle x_{i}x_{j}\rangle$ and the distribution of the variance $\sigma^{2}$ can be extracted for different samples of initial operators. Finally, these two statistical quantities are computed as the system changes from nonchaotic to chaotic. Independent of the initial distributions, the resulting statistics of $\langle x_{i}x_{j}\rangle$ and of $\sigma^{2}$ approach the normal distribution when the size of matrix becomes large, as discussed in Section 3. For the average correlation matrix $\langle x_{i}x_{j}\rangle$, its different patterns in the nonchaotic and the chaotic regime are shown in Figure 5. For the resulting normal distributions of $\sigma^{2}$, the logarithm of its statistical average $\langle\sigma^{2}\rangle$ and standard deviation are shown in Figure 8. In the nonchaotic case, the resulting distributions of $\sigma^{2}$ are almost overlapping together for all the five initial distributions. In the chaotic case, they split into two groups well-separated from each other: one group consists of GOE, URE and UIM, while the other group consists of GUE and UCP.
@@
The paper is organized as follows. In Section 2, we briefly review the relevant physics of the Krylov operator space and explain the sampling of random initial operators. In Section 3, we qualitatively explain the Wishart distribution of $\langle x_{i}x_{j}\rangle$ and the chi-square distribution of $\sigma^{2}$, which approach the normal distribution when the size of matrix is large enough. In Section 4, we show the detailed results for the Sinai billiard. In Section 5, we discuss the physics of our results and conclude with a discussion of open questions.
@@ 2 Preliminary: Krylov operator space
Obviously, results in the Krylov space will vary when the initial operator $\mathcal{O}$ changes. One way to deal with this dependence on initial operators is to do statistics and this is what we do in this paper. We choose samples of random initial operators from different probability distributions and study the resulting statistics of the Lanczos coefficients. Specifically, we consider random initial operators ($\mathcal{O}_{GOE}$, $\mathcal{O}_{GUE}$ and $\mathcal{O}_{uni}$) from the Gaussian orthogonal ensemble (GOE), the Gaussian unitary ensemble (GUE) and the uniform distribution.
These operators are evaluated in the energy representation. In the case of GOE, the random matrix, $\mathcal{O}_{mn}\coloneqq\left\langle m\right|\mathcal{O}_{GOE}\left|n\right\rangle$, is a real symmetric matrix whose matrix elements satisfy the following joint probability density function mehta1991random :
$\mathcal{O}_{mn}^{(0,1)}\sim U(-\sqrt{3},\sqrt{3})$ (2.10)
@@ 3 Statistical properties of Lanczos coefficients
To be specific, We focus on the statistics of the more interesting correlation matrix  [^2] In statistics, this is the scatter matrix to estimate the covariance matrix of a distribution. Since we are not doing pure mathematics here, we prefer to use the more physical term ’correlation matrix’ Eynard:2015aea . $x_{i}x_{j}$, and the resulting distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$ across a sample. Naively, the resulting statistics of $x_{i}x_{j}$ and $f_{\sigma^{2}}$ depend on the distributions of the initial operator $\mathcal{O}$. However, we find that their statistics are the Wishart distribution and the rescaled chi-square distribution respectively, independent of the distribution of initial operators. As the matrix size of the operator increases, they approach the normal distribution, at which time the behavior of $f_{\sigma^{2}}$ depends only on whether the system is nonchaotic or chaotic: nearly complete overlapping $f_{\sigma^{2}}$ in the nonchaotic case, and two well-separated groups of $f_{\sigma^{2}}$ in the chaotic case.
Before showing the explicit results, we qualitatively analyze the statistical propertities of Lanczos coefficients in the remaining of this section. We firstly explain the statistics in a purely intuitive manner. Then we explain the statistics using the Wishart distribution wishart1928generalised .
@@ 3.2 Distribution of the average correlation matrix $\langle x_{i}x_{j}\rangle$
Note that our average correlation matrix $\langle x_{i}x_{j}\rangle$ is exactly the sample covariance matrix $M$, because $\langle x_{i}x_{j}\rangle=M_{ij}/N$. So it can indeed capture the statistics behind the ensemble average and we can use properties of the Wishart distribution to understand its behavior. The Wishart distribution can be viewed as a multi-variate generalization of the chi-squared distribution as follows: if $K=\Sigma=1$, it reduces to the chi-squared distribution with $N$ degrees of freedom. However, this is exactly the case in this paper. The distributions of initial operators we used (2.8)-(2.10) are univariate, whose $\Sigma$ is the identity matrix. Then it is naturally to expect that the resulting $K$-component random vector $\vec{X}$ is also univariate. So we can refine the previous assumption by expecting that $\vec{X}$ is from a univariate normal distribution $N(0,\sigma_{x})$. In this way, the Wishart distribution satisfied by $\langle x_{i}x_{j}\rangle$ reduces to the chi-square distribution of $\langle x_{i}^{2}\rangle$. In Figure 1(a), we show the histogram of $\langle x_{i}^{2}\rangle$ for the case of GUE when the system is chaotic with $\mathcal{N}_{max}=15$. It is indeed the chi-square distribution. As the size of matrix increases, the distribution would become normal by the central limit theorem. In Figure 1(b), the corresponding histogram of $\langle x_{i}x_{j}\rangle$ with $\mathcal{N}_{max}=50$ is plotted and it is indeed the normal distribution. Similarly properties are there when the system is nonchaotic.
(a) Histogram of $\langle x_{i}^{2}\rangle$, $\mathcal{N}_{max}=15$.
(b) Histogram of $\langle x_{i}x_{j}\rangle$, $\mathcal{N}_{max}=50$.
Figure 1: Distribution of $\left\langle x_{i}x_{j}\right\rangle$ in the chaotic case.
@@ 3.3 Distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$
Qualitatively, we can expect that the second term with $i\neq j$ can be neglected compared with the first term. This is true when the set size $K$ is large. In the detailed computation of this paper, this condition is satisfied and the second term is smaller than the first term by at least one order of magnitude. So we obtain qualitatively
@@
In the previous subsection 3.2, we have discussed the physics that $x_{i}$ are independent random variables from a univariate normal distribution $N(0,\sigma_{x})$. If the standard deviation $\sigma_{x}$ is known, we can scale $x_{i}$ by an overall factor to make its standard deviation being $1$. In this way, the resulting distribution $f_{\sigma^{2}}$ (3.18) is proportional to the chi-square distribution by an overall factor, called the rescaled chi-square distribution here. When $\mathcal{N}_{max}$ increases, it would approach the normal distribution.
Numerical results indeed support this expectation. In Figure 2, we show the histogram of the variance $\sigma^{2}$ for the case of GOE when the system is chaotic. When $\mathcal{N}_{max}=5$, the histogram clearly satisfies the rescaled chi-square distribution with the curve being the fitted distribution. As it increases to $\mathcal{N}_{max}=100$, the histogram approaches the normal distribution and the curve are the fitted normal distribution. Similar properties are there for the other ensembles and for the nonchaotic case, as discussed in Section 4 and Appendix C.
(a) $\mathcal{N}_{max}=5$, rescaled chi-square distribution.
(b) $\mathcal{N}_{max}=100$, normal distribution. Skewness=0.201339, Kurtosis=3.19744
Figure 2: Chi-square distribution of $f_{\sigma^{2}}$ in the chaotic case for GOE.
From this section’s analysis, we see that the amount of random numbers is determined by $\mathcal{N}_{max}$, $K$ and the sample size $N$ of initial operators. If they are large enough, we would expect the resulting distribution to be normal. In the following computation, we use the choice $\mathcal{N}_{max}=100$ and $K=500$ in consistent with Hashimoto:2023swv , and we average over $N=5000$ intial operators. This number turns out to be large enough and the resulting distribution function $f_{\sigma^{2}}$ is the normal distribution.
4 Model and Results
For concrete, we study the Sinai billiard which is a typical two-dimensional system of chaos. The Hamiltonian is defined by
@@ 4 Model and Results
The region $\Omega$ is obtained by cutting out a circle of radius $l$ from an equilateral triangle of side length $L$, as shown in Figure 3. The system is integrable when $l=0$ and chaotic when $0<l<L$. For convenient, the area of $\Omega$ is kept at 1 and the parameter $a\equiv l/L$ is introduced, whose value determines the degree of chaos for the system.
Figure 3: The region $\Omega$ of the potential (4.2) of the billiard. Dirichlet conditions are imposed on the boundaries.
@@
For practical computation in the energy representation, we need to set the value of $\mathcal{N}_{max}$ to truncate the number of energy levels. It is adequate to take a value as long as the resulting distributions of $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$ are normal. When this happens, there are enough randomness in the statistics so that $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$ can distinguish the chaotic and the nonchaotic behaviors. Empirically, $\mathcal{N}_{max}=50$ is enough here. For the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$, we use $\mathcal{N}_{max}=50$ for the purpose of numerical efficiency. For the distribution $f_{\sigma^{2}}$ of the variance, we use $\mathcal{N}_{max}=100$ to keep accordance with Hashimoto:2023swv . Then we pick a random matrix $\mathcal{M}_{\mathcal{O}}$ of size $\mathcal{N}_{max}\times\mathcal{N}_{max}$ from the probability distributions (2.8)-(2.10) for the initial random operator $\mathcal{O}$. After that, we compute the Lanczos coefficient $\left\{b_{n}\right\}$ of $\mathcal{K}_{m}(\mathcal{L},\mathcal{M}_{\mathcal{O}})$ by the algorithm in Section 2. To calculate the variance $\sigma^{2}$ of Lanczos coefficients, we choose the set size to be $K=5\mathcal{N}_{max}$, in consistent with Hashimoto:2023swv . Finally, we repeat this process $N=5000$ times to get the resulting distribution of $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$.
In Figure 4, an example of Lanczos coefficients $b_{n}$ for the case of GUE is shown. There is a clear difference between the chaotic and the nonchaotic case, which can be explored further by the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$. The set of coefficients $\left\{b_{n}\right\}$ between the two red lines, $5\mathcal{N}_{max}\leq n\leq 10\mathcal{N}_{max}$, is selected to compute the variance $\sigma^{2}$. Note that this selection of the set does not qualitatively affect the result of this paper. The effect of selecting the set of $\left\{b_{n}\right\}$ from different algorithm step sizes is discussed in the appendix B.

### v3

@@ Abstract
We study the statistical properties of Lanczos coefficients over an ensemble of random initial operators generating the Krylov space. We propose two statistical quantities that are important in characterizing the complexity: the average correlation matrix $\langle x_{i}x_{j}\rangle$ of Lanczos coefficients and the resulting distribution of the variance of Lanczos coefficients. Their resulting statistics are the Wishart distribution and the (rescaled) chi-square distribution respectively, which are independent of the distributions of initial operators and become the normal distribution in the case of large matrix size. As a numerical example, we use the typical billiard system with an integrability-breaking term and choose samples of random initial operators from given probability distributions (GOE, GUE and the uniform distribution). It agrees with the phenomenological analysis and further interesting behaviors are obtained, which indicates a consistent connection between RMT, Anderson localization and Krylov complexity.
@@ 1 Introduction
The physics of complexity is important for understanding many different phenomena, ranging from condensed matter physics to black hole physics. Especially, the operator complexity captures the spread of an operator under time evolution and is an important indicator in the analysis of chaos. Recently the spread of an operator is explored in the Krylov space Parker:2018yvk , where the dynamics depends on the expansion coefficients of the operator in the Krylov basis and on the Lanczos coefficients generated in the process of constructing this Krylov basis. There the Krylov complexity (K-complexity) is constructed from the expansion coefficients and is proposed as an upper bound on the operator complexity, including the famous bound $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena:2015waa of the growth of the out-of-time-order correlation function (OTOC) 1969JETP28:1200L .
@@
Then it leads to an interesting connection between K-complexity and Random Matrix Theory (RMT). By adding an integrability-breaking term to the XXZ spin chain, the late-time saturation value of K-complexity is computed as the system changes from nonchaotic to chaotic behavior Rabinovici:2022beu . And it approaches the behavior of similar values computed with Hamiltonian drawn from RMT ensembles. Also, three different initial operators generating the Krylov space is randomly selected and compared there. In Craps:2024suj , the K-complexity is also studied by choosing multiple initial operators.
@@
Along this line, an important question arises: What would happen for the Lanczos coefficients if we were to perform statistics of initial operators over a probability distribution. In this paper, we study this possibility by proposing two statistical quantities that have a deep connection with Anderson localization and RMT: the average correlation matrix $\langle x_{i}x_{j}\rangle$ of Lanczos coefficients and the resulting distribution of the variance $\sigma^{2}$. We qualitatively analyze their statistical properties based on RMT and on existing results in Krylov complexity. The average correlation matrix $\langle x_{i}x_{j}\rangle$ satisfies the Wishart distribution and the resulting distribution of $\sigma^{2}$ satisfies a rescaled chi-square distribution. This is independent of the initial operator distributions and approaches the normal distribution as the size of the operator matrix increases. For the purpose of characterizing operator complexity, the behavior of two quantities should be consistent with each other because of their deep connection with Anderson localization and RMT.
As a numerical example, we use two typical billiard systems originating from a quantization of classical chaos: the Sinai billiard BERRY1981163 and the Stadium billiard benettin1978numerical ; McDonald:1979zz ; Casati:1980ytd , equipped with an integrability-breaking term. The numerical results indeed verify our qualitative analysis and indicate the consistency between them, which implies an interesting connection among RMT, Anderson localization and Krylov complexity. Furthermore, for the Sinai billiard varing from nonchaotic to chaotic, these two quantities have distinct patterns that can capture this transition. In the nonchaotic regime the correlation matrix $\langle x_{i}x_{j}\rangle$ has a pattern that locally resembles the two-electron wave function of the Anderson localization Weinmann1995h2eOF , which disappears (or fading away by spreading out uniformly) in the chaotic regime. And the resulting distributions of $\sigma^{2}$ are nearly the same in the nonchaotic case, but split into two well-separated groups in the chaotic case. However, there is no such distinction for the Stadium billiard. In both examples, the behavior of these two quantities are consistent with each other.
@@
For details of the numerical example, it is done in the energy representation and operators become matrices. Each time, a random initial operator $\mathcal{O}$ is sampled out by picking the corresponding random matrix from the following distributions: the GOE, the GUE, the uniform distribution of complex numbers (UCP), the uniform distribution of pure real numbers (URE) and of pure imaginary numbers (UIM). Then the Lanczos coefficients are computed in the Krylov space associated with this operator, which give the correlation matrix $x_{i}x_{j}$ and the variance $\sigma^{2}$ of Lanczos coefficients. Repeating this enough times, the average correlation matrix $\langle x_{i}x_{j}\rangle$ and the distribution of the variance $\sigma^{2}$ can be extracted for different samples of initial operators. Finally, these two statistical quantities are computed as the system changes from nonchaotic to chaotic.
@@
The paper is organized as follows. In Section 2, we briefly review the relevant physics of the Krylov operator space. In Section 3, we qualitatively analyze the statistical properties of $\langle x_{i}x_{j}\rangle$ and $\sigma^{2}$, and obtain the Wishart distribution and the rescaled chi-square distribution respectively. They approach the normal distribution when the size of matrix is large enough. In Section 4, we provide numerical examples for the Sinai billiard and the Stadium billiard. The Wishart distribution and the rescaled chi-square distribution are verified and more interesting behaviors are obtained as the system changes from nonchaotic to chaotic. In Section 5, we conclude with a discussion of open questions.
@@ 2 Preliminary: Krylov operator space
Obviously, results in the Krylov space will vary when the initial operator $\mathcal{O}$ changes. One way to deal with this dependence on initial operators is to do statistics by sampling random initial operators from different probability distributions. We will focus on this approach in the remaining of this paper.
@@ 3 Statistical properties of Lanczos coefficients
To be specific, We focus on the statistics of the more interesting correlation matrix  [^2] In statistics, this is the scatter matrix to estimate the covariance matrix of a distribution. Since we are not doing pure mathematics here, we prefer to use the more physical term ’correlation matrix’ Eynard:2015aea . $x_{i}x_{j}$, and the resulting distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$ across a sample. Naively, the resulting statistics of $x_{i}x_{j}$ and $f_{\sigma^{2}}$ depend on the distributions of the initial operator $\mathcal{O}$. However, we find that their statistics are the Wishart distribution wishart1928generalised and the rescaled chi-square distribution respectively after a qualitatively analysis, which are independent of the distribution of initial operators. As the matrix size of the operator increases, they approach the normal distribution. In the remaining of this section, we firstly explain the statistics in a purely intuitive manner. Then we explain the statistics in a phenomenological manner.
@@ 3.2 Distribution of the average correlation matrix $\langle x_{i}x_{j}\rangle$
Note that our average correlation matrix $\langle x_{i}x_{j}\rangle$ is exactly the sample covariance matrix $M$, because $\langle x_{i}x_{j}\rangle=M_{ij}/N$. So it can indeed capture the statistics behind the ensemble average and we can use properties of the Wishart distribution to understand its behavior. In practical computation, $N$ is the sample size of random initial operators and $K$ is the number of selected Lanczos coefficients for each initial operator.
@@ 3.3 Distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$
Qualitatively, we can expect that the second term with $i\neq j$ can be neglected compared with the first term. This is true when the set size $K$ is large. This condition is satisfied in all practical computation of the variance $\sigma^{2}$ and the second term is smaller than the first term by at least one order of magnitude. So we obtain qualitatively
@@
This is the general result and it is hard to tell which distribution it satisfies because $x_{i}$ belongs to a vector $\vec{X}$ from a multivariate normal distribution $N_{K}(0,\Sigma)$. But in the simple case when $x_{i}$ are independent random variables from a univariate normal distribution $N(0,\sigma_{x})$, we can scale $x_{i}$ by an overall factor to make its standard deviation being $1$ and the resulting distribution $f_{\sigma^{2}}$ of the variance (3.18) is proportional to the chi-square distribution by an overall factor. We will provide numerical examples for this simplified case of rescaled chi-square distribution.
For the purpose of practical computation, the amount of random numbers is determined by $\mathcal{N}_{max}$, $K$ and the sample size $N$ of initial operators. If they are large enough, we would expect the resulting distribution to be normal. Usually the choice of $N=5000$, $\mathcal{N}_{max}=100$ and $K=500$ turns out to be large enough for the resulting distribution to be the normal distribution and this choice in consistent with Hashimoto:2023swv .
4 Numerical examples
As an example, we study the Sinai billiard and the Stadium billiard which are typical two-dimensional systems of chaos. The Hamiltonian is defined by
@@ 4 Numerical examples
For the Sinai billiard, the region $\Omega$ is obtained by cutting out a circle of radius $l$ from an equilateral triangle of side length $L$, as shown in Figure 1(a). The system is integrable when $l=0$ and chaotic when $0<l<L$. For the Stadium billiard, the region $\Omega$ is obtained by combining a rectangle and a quarter circle as shown in Figure 1(b). For convenience, the area $\Omega$ is kept at 1 and the parameter $a\equiv l/L$ is introduced, whose value determines the degree of chaos for the system.
(a) The Sinai billiard. Dirichlet conditions are imposed on the boundaries.
(b) The Stadium billiard. Dirichlet conditions are imposed on the boundaries.
Figure 1: Region $\Omega$ of $V(x,y)$ (4.2) for the Sinai billiard and the Stadium billiard.
@@
For practical computation in the energy representation, we need to set the value of $\mathcal{N}_{max}$ to truncate the number of energy levels. It is adequate to take a value as long as the resulting distributions of $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$ are normal. When this happens, there are enough randomness in the statistics so that $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$ can capture the relevant physics. Empirically, $\mathcal{N}_{max}=50$ is enough here. For the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$, we use $\mathcal{N}_{max}=50$ for the purpose of numerical efficiency. For the distribution $f_{\sigma^{2}}$ of the variance, we use $\mathcal{N}_{max}=100$ to keep accordance with Hashimoto:2023swv . For random initial operators, we choose several different probability distributions and compare their effects. Specifically, we consider random initial operators ($\mathcal{O}_{GOE}$, $\mathcal{O}_{GUE}$ and $\mathcal{O}_{uni}$) from the Gaussian orthogonal ensemble (GOE), the Gaussian unitary ensemble (GUE) and the uniform distribution. In the case of GOE, the random matrix, $\mathcal{O}_{mn}\coloneqq\left\langle m\right|\mathcal{O}_{GOE}\left|n\right\rangle$, is a real symmetric matrix whose matrix elements satisfy the following joint probability density function mehta1991random :
$\mathcal{O}_{mn}^{(0,1)}\sim U(-\sqrt{3},\sqrt{3})$ (4.5)
Then we pick a random matrix $\mathcal{M}_{\mathcal{O}}$ of size $\mathcal{N}_{max}\times\mathcal{N}_{max}$ from the above probability distributions (4.3)-(4.5) for the initial random operator $\mathcal{O}$. After that, we compute the Lanczos coefficient $\left\{b_{n}\right\}$ of $\mathcal{K}_{m}(\mathcal{L},\mathcal{M}_{\mathcal{O}})$ by the algorithm in Section 2. To calculate the variance $\sigma^{2}$ of Lanczos coefficients, we choose the set size to be $K=5\mathcal{N}_{max}$, in consistent with Hashimoto:2023swv . Finally, we repeat this process $N=5000$ times to get the resulting distribution of $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$. To keep focus on the main physics, we put numerical details that are not essential to the main discussion into the Appendix.
4.1 Distributions of $\langle x_{i}x_{j}\rangle$ and $\sigma^{2}$
Firstly let’s look at the Wishart distribution $W_{K}(\Sigma,N)$ of the average correlation matrix $\langle x_{i}x_{j}\rangle$. The Wishart distribution can be viewed as a multi-variate generalization of the chi-squared distribution as follows: if $K=\Sigma=1$, it reduces to the chi-squared distribution with $N$ degrees of freedom. However, this is exactly the case in the numerical example: the distributions of initial operators (4.3)-(4.5) are univariate, whose $\Sigma$ is the identity matrix. So it is natural to expect that the resulting $K$-component random vector $\vec{X}$ is also univariate, that is, the multivariate normal distribution $N_{K}(0,\Sigma)$ of $\vec{X}$ reduces to a univariate normal distribution $N(0,\sigma_{x})$. In this way, the Wishart distribution satisfied by $\langle x_{i}x_{j}\rangle$ reduces to the chi-square distribution of $\langle x_{i}^{2}\rangle$. For a small matrix-size with $\mathcal{N}_{max}=15$ (that is away from the central limit), the histogram of $\langle x_{i}^{2}\rangle$ is shown in Figure 2(a) for the Sinai billiard and in Figure 2(c) for the Stadium billiard, which are indeed the chi-square distribution. As the size of matrix increases, the distribution would become normal by the central limit theorem. For $\mathcal{N}_{max}=50$, the corresponding histogram of $\langle x_{i}x_{j}\rangle$ is plotted in Figure 2(b) for the Sinai billiard and in Figure 2(d) for the Stadium billiard, and they are indeed the normal distribution. Here we only show the case when the system is chaotic and the initial distribution is GUE. For other cases there are similar properties, which are omitted because they do not provide extra insights and would ruin the conciseness and readability of the paper.
(a) Sinai billiard, $\langle x_{i}^{2}\rangle$, $\mathcal{N}_{max}=15$.
(b) Sinai billiard, $\langle x_{i}x_{j}\rangle$, $\mathcal{N}_{max}=50$.
(c) Stadium billiard, $\langle x_{i}^{2}\rangle$, $\mathcal{N}_{max}=15$.
(d) Stadium billiard, $\langle x_{i}x_{j}\rangle$, $\mathcal{N}_{max}=50$.
Figure 2: Wishart distribution of $\left\langle x_{i}x_{j}\right\rangle$ in the chaotic case for GUE.
Now let’s look at the resulting distribution $f_{\sigma^{2}}$ of the variance. In this numerical example $x_{i}$ are essentially independent random variables from $N(0,\sigma_{x})$, so the resulting distribution $f_{\sigma^{2}}$ is the rescaled chi-square distribution according to  (3.18). Numerical results indeed support this expectation. For a very small matrix-size with $\mathcal{N}_{max}=5$ (far away from the central limit), the histogram of the variance $\sigma^{2}$ is shown in Figure 3(a) for the Sinai billiard and in Figure 3(c) for the Stadium billiard, which are indeed the rescaled chi-square distribution with the curve being the fitted distribution. As the size of matrix increases, the distribution would become normal by the central limit theorem. For $\mathcal{N}_{max}=100$, the corresponding histogram of the variance $\sigma^{2}$ approaches the normal distribution and is plotted in Figure 3(b) for the Sinai billiard and in Figure 3(d) for the Stadium billiard, where the curve are the fitted normal distribution. Here we only show for the case of GOE when the system is chaotic. For other cases there are similar results, which are omitted again for the conciseness and readability of the paper.
(a) Sinai billiard, $\mathcal{N}_{max}=5$, rescaled chi-square distribution.
(b) Sinai billiard, $\mathcal{N}_{max}=100$, normal distribution. Skewness=0.201339, Kurtosis=3.19744
(c) Stadium billiard, $\mathcal{N}_{max}=5$, rescaled chi-square distribution.
(d) Stadium billiard, $\mathcal{N}_{max}=100$, normal distribution. Skewness=0.18738, Kurtosis=3.05751.
Figure 3: Rescaled chi-square distribution of $f_{\sigma^{2}}$ in the chaotic case for GOE.
4.2 Behaviors between the nonchaotic and the chaotic case
(a) Sinai billiard, $\left|\langle x_{i}x_{j}\rangle\right|$, $a=1$.
(b) Sinai billiard, $\left|\langle x_{i}x_{j}\rangle\right|$, $a=0$.
(c) Stadium billiard, $\left|\langle x_{i}x_{j}\rangle\right|$. $a=1$.
(d) Stadium billiard, $\left|\langle x_{i}x_{j}\rangle\right|$. $a=0$.
Figure 4: Patterns of $\left|\langle x_{i}x_{j}\rangle\right|$ for the Sinai billiard and the Stadium billiard, $\mathcal{N}_{max}=50$.
(a) Logarithm plot for the average $\langle\sigma^{2}\rangle$ of $f_{\sigma^{2}}$. The error bar is the standard deviation. $\mathcal{N}_{max}=100$.
(b) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$, $a=0$, $\mathcal{N}_{max}=50$.
(c) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ , $a=1$, $\mathcal{N}_{max}=50$.
Figure 5: Sinai billiard: behaviors of $f_{\sigma^{2}}$ and $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ when the system changes from nonchaotic ($a=0$) to chaotic ($a=1$).
(a) Plot for the average $\langle\sigma^{2}\rangle$ of $f_{\sigma^{2}}$. The error bar is the standard deviation. $\mathcal{N}_{max}=100$.
(b) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$, $a=0$, $\mathcal{N}_{max}=50$.
(c) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ , $a=1$, $\mathcal{N}_{max}=50$.
Figure 6: Stadium billiard: behaviors of $f_{\sigma^{2}}$ and $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ when the system changes from nonchaotic ($a=0$) to chaotic ($a=1$).